A sum of money invested at 6percent for a certain period of time yield rpt 960 as interst if this sum had been invested for 5 years more it would have yield rpt 2160 as interst find the sum
Answer by addingup(3677)

Invested at simple interest:
x * 0.06 * y = 960
x * 0.06 * y + 5 = 2160
x * 0.06 * 5 = 1200
x * 0.3 = 1200
x = 1200/0.3 = 4000
The amount invested is rpt4000